OpenStudy (anonymous):

derivative sin^2 (X)

OpenStudy (anonymous):

derivative of sin (x) = cos(x) from here we first use the power rule to get \[2\sin(x)(d(\sin(x)/dx)\] the reason for the derivative of the sin(x) again is because of the chain rule. the final answer is \[2\sin(x)\cos(x)\]

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."